a train traveled route fro city A to city B in 4 hours. If the engineer increased the speed of the train by 30 mph, The trip would only take 3 hours. How far is city A from city B?
Answer by ankor@dixie-net.com(22740)

): a train traveled route fro city A to city B in 4 hours.
If the engineer increased the speed of the train by 30 mph, The trip would only take 3 hours.
How far is city A from city B?
:
let s = the slower speed
then
(s+30) = the faster speed
:
Write a distance equation, dist = time * speed
slow speed dist = fast speed dist
4s = 3(s+30)
4s = 3s + 90
4s - 3s = 90
s = 90 mph
Find the dist
4 * 90 = 360 mi
Check the distance
3 * 120 = 360 mi